OpenStudy (anonymous):

What is the area of a sector with radius 10" and measure of arc equal to 45°?

OpenStudy (dumbcow):

take area of circle times ratio of 45/360

ganeshie8 (ganeshie8):

(45/360) x pi x r x r

OpenStudy (anonymous):

would it be 7.85? or no

OpenStudy (unklerhaukus):

\[A_{\text{circle}}=\pi\times r^2\] \[A_{\text{sector}}=\frac{\theta°}{360°}\times\pi\times r^2\]

OpenStudy (unklerhaukus):

In this example \[r=10\text{ in}\]\[\theta°=45°\]

OpenStudy (unklerhaukus):

7.85 is not correct

OpenStudy (anonymous):

haha i accidentally put a 2piR instead of piR^2 I got 39.25 is that correct?

OpenStudy (unklerhaukus):

yes yes

OpenStudy (anonymous):

ah!! haha YES!! haha success